# GNN-DT: Graph Neural Network Enhanced Decision Transformer for Efficient Optimization in Dynamic Environments

### v1

@@ 5.1 Training Performance
Figure 4: Comparison of smart charging algorithms based on (a) State of Charge, and (b) Action Probability.
5.2 Illustrative Example of EV Charging
With the models trained, we proceed to compare the behavior of the best baseline models trained (DT, Q-DT, GNN-DT) against the heuristic BaU algorithm in an EV charging scenario. Fig. 4a presents the SoC progress for three EVs connected one after the other to a single charger throughout the simulation, while Fig. 4b illustrates the charging and discharging actions of all chargers taken by each algorithm.
In Fig. 4a, the heuristic BaU algorithm consistently overcharges the EVs, often exceeding the desired SoC levels. In contrast, both DT and Q-DT fail to fully satisfy the desired SoC, except for the last EV, resulting in suboptimal performance. Conversely, GNN-DT successfully achieves the desired SoC for all EVs, closely mirroring the behavior of the optimal algorithm. This demonstrates GNN-DT’s ability to precisely control charging actions based on dynamic state information. Fig. 4b provides further insights into the actions taken by each algorithm. The optimal solution primarily employs maximum charging or discharging power, since it knows the future. In comparison, GNN-DT exhibits a more refined approach, modulating charging power within a range of -6 to 11 kW. On the other hand, baseline DT and Q-DT display a narrower range of actions, limiting their ability to optimize the charging schedules and adapt to varying conditions. These results underscore the superior capability of GNN-DT in managing the complexities of EV charging dynamics. For a more detailed analysis of key performance metrics in EV charging scenarios, refer to Appendix A.3.
@@ 5.3 Evaluation of Generalization and Scalability
Evaluating the generalization of RL models across varying state transition probabilities is crucial for ensuring consistent performance under diverse conditions (Wang et al., 2020). In Fig. 5a, the generalization capabilities of GNN-DT and other baselines are assessed in environments with small, medium, and extreme variations in state transition probabilities (compared to the training environment). While the baseline methods experience significant performance drops as the evaluation environment deviates from the training setting, GNN-DT maintains strong performance across all scenarios. This highlights the critical role of GNN-based embeddings in improving model robustness and generalization.
A key advantage of the GNN-DT architecture, not present in classic DTs, is its invariance to problem size, i.e. the same network can be applied to both smaller and larger-scale environments. Fig. 5b illustrates the scalability and generalization performance of GNN-DT compared to the BaU algorithm and Optimal policy. GNN-DT, originally trained on a 25-charger environment, is evaluated in environments with 5, 50, 75, and 100 chargers. As expected, performance decreases as the number of chargers increases, since GNN-DT was not trained on larger-scale environments. Nevertheless, it still outperforms the heuristic BaU, demonstrating the model’s capability to handle problem size variation. In future work, training GNN-DT on a range of charger configurations simultaneously could further enhance its adaptability across a broader spectrum of environments.
@@
Table 4: Maximum reward achieved by GNN-DT in a large-scale EV charging optimization task involving 250 chargers.
Total
Trajectories
Avg. Dataset
Reward
GNN-DT
Reward
@@
The scalability and effectiveness of GNN-DT were tested when trained on a significantly larger optimization problem involving 250 charging stations (CSs). In this scenario, the model must handle up to 250 action variables per step and over 1,000 state variables, which include critical information such as power limits and battery levels. The results presented in Table 4 demonstrate that GNN-DT shows promise for addressing more complex optimization tasks. However, the model requires a substantial increase in both the number of training trajectories and memory resources to maintain efficiency, highlighting a well-known limitation of DT-based approaches.
@@ 6 Conclusions
In this work, we introduced a novel DT-based architecture, GNN-DT, which incorporates GNN embedders to significantly enhance sample efficiency and overall performance. Through extensive evaluation across various datasets, including expert, random, and BaU, we demonstrated that traditional DTs and online RL algorithms fail to generalize effectively in real-world settings without specialized embeddings. We further show that both the size and quality of input trajectories critically impact the training process, underscoring the importance of carefully selecting datasets for effective learning. Finally, by leveraging the power of GNN embeddings, GNN-DT improved the model’s ability to generalize in previously unseen environments and handle large, complex action spaces. These contributions demonstrate GNN-DT’s potential to address complex dynamic optimization challenges beyond EV charging. They also underscore the critical roles of data quality and model architecture in enabling efficient real-world deployment.
Acknowledgements
The study was partially funded by the DriVe2X research and innovation project from the European Commission with grant numbers 101056934. The authors acknowledge the use of computational resources of the DelftBlue supercomputer, provided by Delft High Performance Computing Centre (https://www.tudelft.nl/dhpc). This work used the Dutch national e-infrastructure with the support of the SURF Cooperative using grant no. EINF-5716.
@@ Appendix A Appendix: EV Charging Model
The first appendix Section provides the complete mixed integer programming (MIP) formulation of the optimal EV charging problem and presents detailed experimental results for key evaluation metrics.
@@ A.1 Complete EV charging MIP model
The investigated EV charging problem expands over a simulation with $t$ timesteps, $t\in\mathcal{T}$ where a CPO decides the charging and discharging power ($p^{+}_{i,t}$ and $p^{-}_{i,t}$) for every charging station $i\in\mathcal{I}$. Since the chargers can be spread around the city, there are charger groups $w\in\mathcal{W}$, that can have a lower-level aggregated power limits representing connections to local power transformers. Based on these factors, the overall optimization objective is defined as follows:
@@
$e_{j,i,t}=e^{a}_{j,i,t}$ $\forall j,\;\forall i,\;\forall t|\;t={t}^{a}_{j,i,t}$ (17)
@@ A.2 Evaluation Metrics
A.3 Complete Experimental Results of EV Charging.
@@ B.1 Training Hyperparameters
Maximum Epochs $250$ (for small scale), $400$ (for large scale)
@@
Memory per CPU (GB) $8$ (for small scale), $16$ or $40$ (for large scale)
@@
B.2 Detailed Training Curves
@@
B.3 Mixed Dataset Training Curves
Fig. 8 presents the learning curves for the mixed datasets approach, comparing performance across various combinations of Optimal, Random, and BaU datasets for both $K=2$ and $K=10$. Fig. 8a illustrates the results for mixed Optimal datasets, where different proportions of the Optimal and Random datasets (e.g., 50% Optimal + 50% Random) are used for training. The performance becomes more unstable as more Random data is included. Interestingly, the performance for all the mixed datasets demonstrates better maximum reward reached compared to the Optimal-only dataset, highlighting the benefits of combining high-quality and lower-quality data. Fig. 8b shows similar trends for the Mixed-BaU datasets. While the BaU dataset alone performs worse than the Optimal dataset, mixing it with Random data still yields improvements, with the 75% BaU and 25% Random combination showing the best results. The results underscore the potential of mixing datasets to improve training performance, especially when high-quality data (Optimal or BaU) is supplemented with lower-quality Random data.

### v2

@@ 5.2 Illustrative Example of EV Charging
Figure 3: Comparison of smart charging algorithms for a single simulation day.
In Fig. 3a, the heuristic BaU algorithm consistently overcharges the EVs, often exceeding the desired SoC levels. In contrast, both DT and Q-DT fail to satisfy the desired SoC. Conversely, GNN-DT successfully achieves the desired SoC for all EVs, closely mirroring the behavior of the optimal algorithm. This demonstrates GNN-DT’s ability to precisely control charging based on dynamic state information. Fig. 3b provides further insights into the actions taken by each algorithm. The optimal solution primarily employs maximum charging or discharging power, since it knows the future. In comparison, GNN-DT exhibits a more refined approach, modulating charging power within a range of -6 to 11 kW. Baseline DT and Q-DT display a narrower range of actions, limiting their ability to optimize the charging schedules and adapt to varying conditions. These results underscore the superior capability of GNN-DT in managing the complexities of EV charging dynamics. For a more detailed analysis of key performance metrics in EV charging scenarios, refer to Appendix A.4.
@@ 5.3 Evaluation of Generalization and Scalability
Evaluating the generalization of RL models across varying state transition probabilities is crucial for ensuring consistent performance under diverse conditions [40]. In Fig. 4a, the generalization capabilities of GNN-DT and other baselines are assessed in environments with small, medium, and extreme variations in state transition probabilities (compared to the training environment, see Appendix A.5 for more details about the different distributions). While the baseline methods experience significant performance drops as the evaluation environment deviates from the training setting, GNN-DT maintains strong performance across all scenarios. This highlights the critical role of GNN-based embeddings in improving model robustness and generalization. A key advantage of the GNN-DT architecture, not present in classic DTs, is its invariance to problem size, i.e., the same RL agent can be applied to both smaller and larger-scale environments. Fig. 4b illustrates the scalability and generalization performance of GNN-DT compared to the BaU algorithm and Optimal policy. GNN-DT, trained on a 25-charger setup, was tested on 5, 50, 75, and 100-charger environments. While its performance predictably declines at larger scales, since it wasn’t trained on them, it still outperforms the BaU heuristic, demonstrating robustness to problem-size variation. Training GNN-DT on a mix of charger numbers could potentially further improve its adaptability.
@@
The scalability and effectiveness of GNN-DT were tested when trained on a significantly larger optimization problem involving 250 charging stations. In this scenario, the model must handle up to 250 action variables per step and over 1,000 state variables, which include critical information such as power limits and battery levels. The results presented in Table 4 demonstrate that GNN-DT shows promise for addressing more complex optimization tasks. However, the model requires a substantial increase in both the number of training trajectories and memory resources to maintain efficiency, highlighting a well-known limitation of DT-based approaches. Scaling the problem 10× roughly multiplies GPU memory usage, e.g. storing 3,000 trajectories takes $\approx 2$ GB for 25 chargers versus $\approx 20$ GB for 250. While this can bottleneck large-scale training, parallelization and mini-batching mitigate it, and overall compute scales with the transformer’s context length K (see Fig. 8), pointing to interesting directions for very large problem graphs as future work.
Table 4: Max. reward of GNN-DT in a large-scale EV charging optimization task with 250 chargers.
Total Trajectories
Avg. Dataset Reward
GNN-DT Reward
@@ 6 Conclusions
In this work, we introduced a novel DT-based architecture, GNN-DT, which incorporates GNN embedders to significantly enhance sample efficiency and overall performance in the sequential decision-making problem of EV charging. Through extensive evaluation across various datasets, including optimal, random, and BaU, we demonstrated that traditional DTs, online and offline RL algorithms fail to effectively solve real-world problems without specialized embeddings. We further show that both the size and quality of input trajectories critically impact the training process, underscoring the importance of carefully selecting datasets for effective learning. Finally, by leveraging the power of GNN embeddings, GNN-DT improved the model’s ability to generalize in previously unseen environments and handle large, complex action spaces. These contributions demonstrate GNN-DT’s potential to address complex dynamic optimization challenges beyond EV charging.
@@ Appendix A Appendix: EV Charging Model
Table 5: Notation for the EV Charging Optimization Problem
Symbol Name Description
Sets
$\mathcal{T}$ Set of timesteps Time horizon for optimization
$\mathcal{I}$ Set of charging stations All EV charging stations
$\mathcal{W}$ Set of charger groups Chargers grouped by local transformer connections
$\mathcal{J}_{i}$ Set of charging sessions Charging sessions at charger $i$
Indexes
$t$ Timestep index Discrete time step
$i$ Charger index Individual EV charging station
$j$ Session index Charging session at a charger
$w$ Charger group index Charger groups connected to transformers
Parameters
$t^{a}_{j,i}$ Arrival time Time EV $j$ arrives at charger $i$
$t^{d}_{j,i}$ Departure time Time EV $j$ departs charger $i$
$e^{*}_{j,i}$ Desired battery capacity Desired battery energy at departure for session $j$ at charger $i$
$e^{a}_{j,i}$ Arrival battery energy Battery energy at EV arrival
$\underline{e}_{j,i},\overline{e}_{j,i}$ Battery limits Min/max allowable battery energy
$\underline{p}^{+}_{j,i},\overline{p}^{+}_{j,i}$ Charging power limits Min/max charging power
$\underline{p}^{-}_{j,i},\overline{p}^{-}_{j,i}$ Discharging power limits Min/max discharging power
$p^{*}_{t}$ Total power limit Desired aggregated power
$\Pi^{+}_{t},\Pi^{-}_{t}$ Electricity prices Prices for charging/discharging at timestep $t$
$\Delta t$ Time interval Duration of each timestep
$\overline{p}_{w,t}$ Group power limit Power limit for group $w$ at timestep $t$
Variables
$p^{+}_{i,t},p^{-}_{i,t}$ Charging/discharging power Power assigned at charger $i$, timestep $t$
$\omega^{+}_{i,t},\omega^{-}_{i,t}$ Binary operation indicators Indicates if charger $i$ charges ($+$) or discharges (--) at timestep $t$
$e_{j,i,t}$ EV battery energy Battery level for session $j$ at charger $i$, timestep $t$
$p^{\sum}_{t}$ Total aggregated power Net total power across all chargers at timestep $t$
The first appendix Section provides the complete mixed integer programming (MIP) and MDP formulation of the optimal EV charging problem and presents detailed experimental results for key evaluation metrics.
@@ A.1 Complete EV charging MIP model
The investigated EV charging problem (See notation Table 5) expands over a simulation with $t$ timesteps, $t\in\mathcal{T}$ where a CPO decides the charging and discharging power ($p^{+}_{i,t}$ and $p^{-}_{i,t}$) for every charging station $i\in\mathcal{I}$. Since the chargers can be spread around the city, there are charger groups $w\in\mathcal{W}$, that can have a lower-level aggregated power limits representing connections to local power transformers. Based on these factors, the overall optimization objective is defined as follows:
@@
$e_{j,i,t}=e^{a}_{j,i}$ $\forall j,\;\forall i,\;\forall t|\;t={t}^{a}_{j,i}$ (14)
@@ A.2 Evaluation Metrics
A.3 The MDP State Space
Figure 5: The EV charging problem graph structure and state. (a) shows the raw physical layout of EVs, chargers and transformer with information and power links; (b) recasts this as a four-type node graph (EV, charger, transformer, CPO) carrying only the observations available to the CPO; and (c) prunes any charger without connected EV resulting in a compact dynamic heterogeneous state graph.
At each time step $t$, the environment state $\boldsymbol{s}_{t}\in\mathcal{S}$ is encoded as a heterogeneous graph $\mathcal{G}_{t}=(\mathcal{N}_{t},\mathcal{E}_{t})$, where $\mathcal{N}_{t}$ is the set of nodes (EVs, chargers, transformer, and CPO) and $\mathcal{E}_{t}$ the set of edges encoding physical and informational connections. Each node $n\in\mathcal{N}_{t}$ carries a feature vector $\boldsymbol{x}^{t}_{n}\in\mathbb{R}^{d}$, encapsulating all observations available to the CPO under V2G protocols. Fig. 5 shows how the EV charging problem is modelled as a graph with each node type having different number of features as detailed below.
An EV node’s feature vector:
$\boldsymbol{x}^{\mathrm{ev}}_{t}=\bigl{[}\text{SoC}_{t},\;t^{\rm d}-t,\;j,\;i,\;w\bigr{]}$ (23)
encodes its current state of charge $\text{SoC}_{t}\in[0,1]$, the remaining steps until departure $t^{\rm d}-t$, and the identifiers $j$, $i$, and $w$ of the charging point, charger, and transformer group to which it is connected. Exact SOC and departure-time information are assumed available via ISO 15118-style V2G signaling. A charging station (cs) node $i$ is described by:
$\boldsymbol{x}^{\mathrm{cs}}_{t}=\bigl{[}\overline{p}^{+}_{i,j},\;\overline{p}^{-}_{i,j},\;i\bigr{]},$ (24)
where $\overline{p}^{+}$ and $\overline{p}^{-}$ are its maximum charging and discharging power and $i$ its unique identifier. A transformer (tr) node $w$ has a feature vector:
$\boldsymbol{x}^{\mathrm{tr}}_{t}=\bigl{[}\overline{p}_{w,t},\;w\bigr{]},$ (25)
where $\overline{p}_{w,t}$ denotes its maximum power capacity and $w$ its identifier, ensuring the policy respects feeder-level constraints. The CPO (system) node aggregates temporal and economic context in
$\boldsymbol{x}^{\mathrm{cpo}}_{t}=\Bigl{[}\tfrac{d}{7},\;\sin\!\bigl{(}\tfrac{h}{48\pi}\bigr{)},\;\cos\!\bigl{(}\tfrac{h}{48\pi}\bigr{)},\;\Pi^{+}_{t},\;p^{\sum}_{t-1}\Bigr{]}.$ (26)
Here $d$ and $h$ encode day-of-week and hour to capture demand patterns, $\Pi^{+}_{t}$ is the charging price at $t$, and $p^{\sum}_{t-1}$ is the previous time step’s total power draw, providing feedback for profit-maximizing decisions.
A.4 Complete Experimental Results of EV Charging.
@@ A.3 The MDP State Space
A.5 State Transition Distribution Experiments
To analyze the generalization capability of our model, we examine how it performs under environments with varying state transition probabilities. Specifically, we visualize key environment variables that directly impact the transition dynamics. Figure 7a–d presents the probability distributions of EV arrival time, departure time, duration of stay, and state of SoC at arrival across four scenarios: the original training environment and environments with small, medium, and extreme variations. These plots help quantify the extent of variation in each case. Additionally, Figure 7e illustrates the temporal distribution of the power limit in each scenario, providing further insight into the differences in environment configuration.
(a) Arrival time
(b) Departure time
(c) Time of stay
(d) SoC at arrival
(e) Power limit
Figure 7: Overview of the five key metrics: (a) arrival time, (b) departure time, (c) power limit, (d) state of charge upon arrival, and (e) time of stay.
@@ B.1 Training Hyperparameters
Number of Epochs $250$ (for small scale), $400$ (for large scale)
@@
CPU Memory (GB) $8$ (for small scale), $16$ or $40$ (for large scale)
@@
B.2 Impact of larger context lengths (K)
Figure 8: GNN-DT performance for larger context lengths (K).
B.3 Detailed Training Curves
@@
B.4 Mixed Dataset Training Curves
Fig. 10 presents the learning curves for the mixed datasets approach, comparing performance across various combinations of Optimal, Random, and BaU datasets for both $K=2$ and $K=10$. Fig. 10a illustrates the results for mixed Optimal datasets, where different proportions of the Optimal and Random datasets (e.g., 50% Optimal + 50% Random) are used for training. The performance becomes more unstable as more Random data is included. Interestingly, the performance for all the mixed datasets demonstrates better maximum reward reached compared to the Optimal-only dataset, highlighting the benefits of combining high-quality and lower-quality data. Fig. 10b and Table 8 shows similar trends for the Mixed-BaU datasets. While the BaU dataset alone performs worse than the Optimal dataset, mixing it with Random data still yields improvements, with the 75% BaU and 25% Random combination showing the best results. The results underscore the potential of mixing datasets to improve training performance, especially when high-quality data (Optimal or BaU) is supplemented with lower-quality Random data.
@@ B.4 Mixed Dataset Training Curves
Table 8: Maximum reward of GNN-DT trained on merged BaU-Random datasets for $K=2$ and $K=10$.The bold indicates the training dataset with the highest evaluation reward.
Dataset Total Traj. Avg. Dataset Reward GNN-DT Reward (×$\times 10^{5}$)
K=2 k=10
Random (Rnd.) 100% $1000$ $-2.37$$0.39$ $-0.863$ $-1.187$
BaU 25% + Rnd. 75% $1000$ $-1.93$$0.80$ $-0.578$ $-0.461$
BaU 50% + Rnd. 50% $1000$ $-1.51$$0.87$ $-0.665$ $\mathbf{-0.447}$
BaU 75% + Rnd. 25% $1000$ $-1.09$$0.76$ $\mathbf{-0.421}$ $-0.471$
BaU 100% $1000$ $-0.01$$0.01$ $-0.654$ $-0.572$